OpenStudy (sweetemy):

function question: Given h(x)=5, find: a. h(2) b. h(x+3) c. h(-x)

OpenStudy (jsblumen):

Do you know where to start with this?

OpenStudy (sweetemy):

No

OpenStudy (jsblumen):

so for a function, like f(x)=x, if it's f(1)=1, you just replace the x with your number. In this case, you may be overthinking it.

OpenStudy (sweetemy):

How am I overthinking it?

OpenStudy (sweetemy):

how can you solve h(2) if there is nothing to plug in? I'm just really confused

OpenStudy (jsblumen):

Right, so what does the function tell you? No matter what you plug in for x, the answer will be 5.

OpenStudy (sweetemy):

So the answer is h(2)=5?

OpenStudy (jsblumen):

Yep!

OpenStudy (sweetemy):

Okay thank you, so the second one is just h(2+3)=5?

OpenStudy (sweetemy):

Also I'm a bit confused why h(2) would equal five, like logically speaking. I understand the concept but still.

OpenStudy (jsblumen):

If you graph the function, it would be a horizontal line at y=5, so no matter what x equals, the answer will always be five.

OpenStudy (sweetemy):

That makes sense, thank you!

OpenStudy (sweetemy):

would the second one be h(2+3)=5 and the third one h(-5)=5?

OpenStudy (stamp):

Examples to clarify: if h(x)=5+x, then h(1)= 5 +1 and h(1)=6 if h(x)=5x, then h(1)= 5(1) and h(1)=5 if h(x)=5, then h(1)= 5, and h(1)=5

OpenStudy (sweetemy):

Thank you this is easier to understand.